# Dual Ascent Diffusion for Inverse Problems

### v1

@@ 4.1 Experimental Setup
We evaluate our method on two image datasets, FFHQ $256\times 256$ [19] and ImageNet $256\times 256$ [10]. We utilize pretrained models from [7] on the FFHQ dataset and from [11] on the ImageNet dataset. We randomly selected 100 images from the validation set for both datasets, and the images were normalized to $[-1,1]$. Our main evaluation metrics include peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), learned perceptual image patch similarity (LPIPS) [41], and residual error, defined as $\|\mathbf{y}-\mathcal{A}(\mathbf{x})\|_{2}^{2}-\sigma^{2}$ (giving zero in expectation for an ideal solver).
@@
Our method is evaluated on multiple inverse problems. Linear tasks include super-resolution (4×4$\times$4 × downsampling), Gaussian deblurring, motion deblurring, inpainting with a $128\times 128$ box, and inpainting with a random mask that removes $70\%$ of the pixels. Nonlinear tasks consist of phase retrieval (oversampling ratio of $2.0$), for which we report the best result out of five runs due to the intrinsic instability of the task, nonlinear deblurring, and high dynamic range (2×2$\times$2 × dynamic range). All measurements include additive Gaussian noise ($\sigma=0.05$).
@@
We compare against the state of the art: DAPS [39], DPS [7], and DiffPIR [42]. Notably, DiffPIR was not proven to handle nonlinear tasks. We closely evaluate our method against DAPS as it achieves the best results among the baselines.
4.2 Main Results
Quantitative evaluation for the linear and nonlinear tasks on FFHQ and ImageNet datasets are shown in Table 1. Our method outperforms the baselines on the vast majority of the tasks, especially in terms of perceptual similarity and residual error. This is further demonstrated in Fig. 1, where we show a qualitative comparison between the baselines and our method. Overall, DDiff reconstructs finer details with less visual artifacts. More examples of each task and choice of hyperparameters are shown in the supplements.
Moreover, DDiff exhibits significantly increased robustness to higher measurement noise. As shown in Fig. 3, although PSNR may be slightly lower than DAPS in the very low measurement noise region (approximately $\sigma<0.05$), PSNR degrades less rapidly as $\sigma$ increases. LPIPS is generally lower for DDiff than that of DAPS in all levels of noise. LPIPS follows a similar trend, especially on the phase retrieval task, where DDiff divides LPIPS by 3 in the high-noise region ($\sigma=0.3$). To visualize the reconstruction quality of DDiff compared to DAPS at the extreme case of measurement noise level ($\sigma=0.3$), refer to Fig. 4. These results suggest that DDiff could be more appropriate for solving challenging inverse problems where measurement noise is prominent, such as but not limited to low-dose CT reconstruction, cryo-electron microscopy, and sparse-view tomography.
Table 1: Quantitative evaluation. Comparing different methods for 5 linear and 3 nonlinear tasks on FFHQ and ImageNet datasets. This evaluation uses 100 validation images and reports the average metric value. The best and second-best results are distinguished by bold and underlined marks, respectively. All tasks are run with a noise of standard deviation $\sigma=0.05$.
FFHQ ImageNet Task Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) Super Resolution 4× DDiff (ours) 30.07 0.824 0.211 0.0028 25.81 0.656 0.396 0.0038 DAPS 29.34 0.783 0.190 0.0029 25.44 0.636 0.295 0.2871 DPS 24.42 0.486 0.346 0.0050 21.10 0.351 0.408 0.0052 DiffPIR 23.71 0.440 0.423 0.0087 20.75 0.312 0.517 0.0093 Inpainting (Box) DDiff (ours) 24.88 0.831 0.110 0.0077 21.15 0.743 0.240 0.0119 DAPS 24.52 0.742 0.174 0.0099 21.22 0.714 0.230 0.2050 DPS 24.68 0.810 0.079 0.0033 19.63 0.725 0.254 0.0412 DiffPIR 19.02 0.527 0.252 0.0106 16.02 0.520 0.329 0.0132 Inpainting (Random) DDiff (ours) 33.08 0.877 0.050 0.0205 28.39 0.758 0.136 0.0241 DAPS 30.76 0.801 0.156 0.0293 27.32 0.725 0.189 0.0788 DPS 30.79 0.807 0.083 0.0217 27.31 0.737 0.235 0.0980 DiffPIR 18.53 0.362 0.622 0.0264 15.82 0.191 0.842 0.1080 Gaussian Deblurring DDiff (ours) 28.87 0.800 0.119 0.0026 22.29 0.471 0.415 0.0046 DAPS 29.63 0.789 0.177 0.0027 25.90 0.658 0.269 0.2373 DPS 27.77 0.704 0.140 0.0029 23.07 0.628 0.332 0.0084 DiffPIR 26.16 0.624 0.297 0.0031 21.64 0.393 0.497 0.0093 Motion Deblurring DDiff (ours) 28.24 0.785 0.129 0.0058 24.16 0.585 0.242 0.0079 DAPS 29.17 0.797 0.186 0.0059 26.61 0.710 0.241 0.1825 DPS 27.93 0.714 0.130 0.0061 23.36 0.611 0.321 0.0082 DiffPIR 22.01 0.327 0.499 0.0074 18.93 0.248 0.586 0.0084 Phase Retrieval DDiff (ours) 29.94 0.816 0.120 1.4690 18.54 0.494 0.262 1.4659 DAPS 29.60 0.768 0.182 1.4770 20.23 0.449 0.397 1.4685 DPS 22.24 0.540 0.307 1.5412 16.03 0.396 0.444 1.5540 DiffPIR 10.04 0.036 0.783 1.6711 9.61 0.021 0.794 1.5410 Nonlinear Deblurring DDiff (ours) 31.48 0.873 0.120 0.0027 29.68 0.805 0.207 0.0035 DAPS 28.45 0.764 0.188 0.0042 27.28 0.718 0.213 0.2643 DPS 25.39 0.643 0.258 0.0095 28.42 0.691 0.271 0.0053 DiffPIR 19.79 0.331 0.583 0.0273 22.13 0.459 0.435 0.0167 High Dynamic Range DDiff (ours) 26.05 0.873 0.129 0.0459 26.50 0.800 0.108 0.0541 DAPS 27.39 0.846 0.163 0.0505 26.10 0.825 0.171 0.7172 DPS 25.79 0.793 0.165 0.0734 22.72 0.721 0.273 0.4951 DiffPIR 17.69 0.645 0.296 0.1292 18.23 0.637 0.289 0.1105
Figure 2: Qualitative results. DDiff demonstrates sharper and cleaner results compared to DPS and DAPS. All tasks are run with a noise of standard deviation $\sigma=0.05$.
@@ 4.2 Main Results
Figure 4: Qualitative results at high measurement noise level. We use 10 randomly-chosen validation images and show the results of phase retrieval at $\sigma=0.3$, comparing our method to the state-of-the-art algorithm DAPS [39].
4.3 Ablation Studies
Evaluation of time efficiency and quality of samples.
One of the most limiting factors of sampling speed for diffusion-based methods is the number of neural function evaluations (NFEs). It measures the number of times the underlying neural network is called during computation, which in our case is the number of times we call the score model during inference. Therefore, we assess the average sampling time and sample quality of our method and of existing baselines as a function of the number of NFEs in Fig. 5. For the same number of NFEs, DDiff achieves better perceptual quality and faster sampling speed compared to DAPS. The enhanced computational efficiency is a consequence of two key factors: reduced backpropagation requirements for gradient calculations and the absence of supplementary MCMC procedures within sampling iterations, both of which are present in the DAPS framework.
@@ 4.3 Ablation Studies
Table 2: Ablating the dual variable and noising step. We compare DDiff (with noise, with $\mathbf{u}$) to other variants, including Diff-PnP-HQS (no noise, no $\mathbf{u}$), Diff-PnP-ADMM (no noise, with $\mathbf{u}$), and DDiff-HQS (with noise, no $\mathbf{u}$). All methods are compared in a fair manner by using the same DDIM noise schedule. This evaluation uses 10 FFHQ validation images and reports the average metric value. The best and second-best results are distinguished by bold and underlined marks, respectively. DDiff, with the dual variable and noising step, significantly outperforms other methods.
5 Discussion

### v2

@@ 4.1 Experimental Setup
We evaluate our method on two image datasets, FFHQ $256\times 256$ [20] and ImageNet $256\times 256$ [10]. For pixel-space diffusion models, we utilize pretrained models from [7] on the FFHQ dataset and from [11] on the ImageNet dataset. For latent diffusion models, we used unconditional LDM-VQ4 (autoencoder with a downsampling factor of 4) pretrained models from [32] for FFHQ and [29] for ImageNet. We randomly selected 100 images from the validation set for both datasets, and the images were normalized to [1,1][-1,1]. Our main evaluation metrics include peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), learned perceptual image patch similarity (LPIPS) [45], and residual error, defined as $\|\mathbf{y}-\mathcal{A}(\mathbf{x})\|_{2}^{2}-\sigma^{2}$ (see Appendix B for details), which quantifies the degree of data consistency.
@@
Our method is evaluated on multiple inverse problems. Linear tasks include super-resolution ($4\times$ downsampling), Gaussian deblurring, motion deblurring, inpainting with a $128\times 128$ box, and inpainting with a random mask that removes $70\%$ of the pixels. Nonlinear tasks consist of phase retrieval (oversampling ratio of $2.0$), for which we report the best result out of five runs due to the intrinsic instability of the task, nonlinear deblurring, and high dynamic range ($2\times$ dynamic range). All measurements include additive Gaussian noise ($\sigma=0.05$).
Figure 2: Qualitative results. DDiff demonstrates sharper and cleaner results compared to DPS [7] and DAPS [43]. All tasks are run with a noise of standard deviation $\sigma=0.05$.
Table 1: Quantitative evaluation. Comparing different methods for 5 linear and 3 nonlinear tasks on FFHQ and ImageNet datasets. This evaluation uses 100 validation images and reports the average metric value. The best and second-best results are distinguished by bold and underlined marks, respectively. All tasks are run with a noise of standard deviation $\sigma=0.05$. See Appendix D for 95% confidence intervals.
FFHQ ImageNet Task Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) Super Resolution 4× DDiff (ours) 30.07 0.824 0.211 0.0028 25.81 0.656 0.396 0.0038 DAPS 29.34 0.783 0.190 0.0029 25.44 0.636 0.295 0.0047 DMPlug 28.55 0.742 0.220 0.0038 24.22 0.649 0.432 0.0039 DPS 24.42 0.486 0.346 0.0050 21.10 0.351 0.408 0.0052 DiffPIR 23.71 0.440 0.423 0.0087 20.75 0.312 0.517 0.0093 DCDP 26.65 0.641 0.410 0.0079 23.51 0.542 0.460 0.0060 DDRM 26.32 0.763 0.286 0.0075 22.26 0.513 0.473 0.0072 Inpainting (Box) DDiff (ours) 24.88 0.831 0.110 0.0077 21.15 0.743 0.240 0.0119 DAPS 24.12 0.742 0.174 0.0099 21.22 0.714 0.230 0.0150 DPS 23.68 0.810 0.079 0.0033 19.63 0.725 0.254 0.0412 DiffPIR 19.02 0.527 0.252 0.0106 16.02 0.520 0.329 0.0145 DCDP 23.67 0.729 0.232 0.0101 20.45 0.732 0.248 0.0132 DDRM 22.15 0.701 0.209 0.0099 18.52 0.713 0.254 0.0194 RED-diff 14.57 0.578 0.586 0.0451 13.98 0.628 0.367 0.0437 Inpainting (Random) DDiff (ours) 33.08 0.877 0.050 0.0205 28.39 0.758 0.136 0.0241 DAPS 30.76 0.801 0.156 0.0293 27.32 0.725 0.189 0.0788 DMPlug 31.65 0.852 0.137 0.0290 26.09 0.740 0.245 0.0316 DPS 30.79 0.807 0.083 0.0217 27.31 0.737 0.235 0.0980 DiffPIR 18.53 0.362 0.622 0.0264 15.82 0.191 0.842 0.1080 DCDP 25.67 0.757 0.224 0.0398 20.40 0.723 0.253 0.1001 Gaussian Deblurring DDiff (ours) 28.87 0.800 0.119 0.0026 22.29 0.471 0.415 0.0046 DAPS 29.63 0.789 0.177 0.0027 25.90 0.658 0.269 0.0084 DMPlug 22.98 0.537 0.288 0.0036 14.82 0.188 0.680 0.0147 DPS 27.77 0.704 0.140 0.0029 21.07 0.528 0.392 0.0073 DiffPIR 26.16 0.624 0.297 0.0031 21.64 0.393 0.497 0.0093 DCDP 16.75 0.173 0.701 0.0141 16.06 0.183 0.674 0.0195 DDRM 24.87 0.725 0.246 0.0052 21.14 0.457 0.464 0.0113 RED-diff 12.18 0.149 1.232 0.0348 12.22 0.128 0.613 0.0256 Motion Deblurring DDiff (ours) 28.24 0.785 0.129 0.0058 24.16 0.585 0.242 0.0079 DAPS 29.17 0.797 0.186 0.0059 26.61 0.710 0.241 0.0085 DMPlug 21.95 0.512 0.304 0.0076 14.81 0.170 0.696 0.0199 DPS 27.93 0.714 0.130 0.0061 23.36 0.611 0.321 0.0082 DiffPIR 22.01 0.327 0.499 0.0074 18.93 0.248 0.586 0.0084 DCDP 9.536 0.039 0.855 0.0547 9.491 0.066 0.771 0.0511 Phase Retrieval DDiff (ours) 29.94 0.816 0.120 0.0040 18.54 0.494 0.262 0.0059 DAPS 29.60 0.768 0.182 0.0049 20.23 0.449 0.397 0.0085 DPS 22.24 0.540 0.307 0.0514 16.03 0.396 0.444 0.1040 DiffPIR 10.04 0.036 0.783 0.1811 9.61 0.021 0.794 0.2410 DCDP 15.20 0.420 0.616 0.1060 11.63 0.201 0.700 0.1220 RED-diff 14.88 0.386 0.656 0.1721 13.89 0.266 0.639 0.1121 Nonlinear Deblurring DDiff (ours) 31.48 0.873 0.120 0.0027 29.68 0.805 0.207 0.0035 DAPS 28.45 0.764 0.188 0.0042 27.28 0.718 0.213 0.0048 DMPlug 27.17 0.791 0.187 0.0051 22.99 0.603 0.366 0.0104 DPS 25.39 0.643 0.258 0.0095 28.42 0.691 0.271 0.0053 DiffPIR 19.79 0.331 0.583 0.0273 22.13 0.459 0.435 0.0167 DCDP 28.87 0.852 0.177 0.0038 25.22 0.700 0.299 0.0105 RED-diff 29.89 0.783 0.185 0.0040 28.07 0.624 0.306 0.0051 High Dynamic Range DDiff (ours) 26.05 0.873 0.129 0.0459 26.50 0.800 0.108 0.0541 DAPS 27.39 0.846 0.163 0.0505 26.10 0.825 0.171 0.0717 DPS 25.79 0.793 0.165 0.0734 22.72 0.721 0.273 0.1951 DiffPIR 17.69 0.645 0.296 0.1292 18.23 0.637 0.289 0.2105 RED-diff 21.28 0.431 0.359 0.0921 21.02 0.567 0.479 0.1322
@@
We compare against the state of the art: DAPS [43], DMPlug [40], DCDP [24], RED-diff [27], DDRM [21], DPS [7], and DiffPIR [46]. Notably, DiffPIR and DDRM were not proven to handle nonlinear tasks. We closely evaluate our method against DAPS as it achieves the best results among the baselines. For latent diffusion adaptation of our method, we compare against LatentDAPS [43], PSLD [30], and ReSample [32]. Note that PSLD cannot handle nonlinear inverse tasks.
@@
Figure 4: Qualitative results at high measurement noise level. We compare our method (DDiff) with DAPS [43] on the phase retrieval task at $\sigma=0.3$, using 10 randomly selected validation images. Despite the severe measurement noise, DDiff successfully recovers coherent global structures and facial semantics, whereas DAPS reconstructions exhibit strong corruption and noise artifacts.
4.2 Main Results
Quantitative evaluation for the linear and nonlinear tasks on FFHQ and ImageNet datasets are shown in Table 1. Our method outperforms the baselines on the vast majority of the tasks, especially in terms of perceptual similarity and residual error. This is further demonstrated in Fig. 2, where we show a qualitative comparison between the baselines and our method. Overall, DDiff reconstructs finer details with fewer visual artifacts. Additional qualitative examples and hyperparameter details are provided in the supplements.
Moreover, DDiff exhibits significantly increased robustness to higher measurement noise. As shown in Fig. 3, although PSNR may be slightly lower than DAPS in the very low measurement noise region (approximately $\sigma<0.05$), PSNR degrades less rapidly as $\sigma$ increases. LPIPS is generally lower for DDiff than that of DAPS in all levels of noise. LPIPS follows a similar trend, especially on the phase retrieval task, where DDiff divides LPIPS by 3 in the high-noise region ($\sigma=0.3$). To visualize the reconstruction quality of DDiff compared to DAPS at the extreme case of measurement noise level ($\sigma=0.3$), refer to Fig. 4. These results suggest that DDiff could be more appropriate for solving challenging inverse problems where measurement noise is prominent, such as low-dose CT reconstruction.
@@ 4.2 Main Results
Table 2: Ablating the dual variable and noising step. We compare DDiff (with noise, with $\mathbf{u}$) to other variants, including Diff-PnP-HQS (no noise, no $\mathbf{u}$), Diff-PnP-ADMM (no noise, with $\mathbf{u}$), and DDiff-HQS (with noise, no $\mathbf{u}$). Full algorithmic descriptions of each variant are provided in Appendix F. All methods are compared in a fair manner by using the same DDIM noise schedule. This evaluation uses 10 FFHQ validation images and reports the average metric value. The best and second-best results are distinguished by bold and underlined marks, respectively. DDiff, with the dual variable and noising step, significantly outperforms other methods.
4.3 Ablation Studies
Evaluation of time efficiency and quality of samples.
One of the most limiting factors of sampling speed for diffusion-based methods is the number of neural function evaluations (NFEs). It measures the number of score model evaluations performed during inference. Therefore, we assess the average sampling time and sample quality of our method and of existing baselines as a function of the number of NFEs in Fig. 5. For the same number of NFEs, DDiff achieves better perceptual quality and faster sampling speed compared to DAPS. The enhanced computational efficiency is a consequence of two key factors: reduced backpropagation requirements for gradient calculations and the absence of supplementary MCMC procedures within sampling iterations, both of which are present in the DAPS framework.
@@ 4.3 Ablation Studies
5 Conclusion